half life of two samples are 0.1 and 0.4 sec  their onitial con are 200 and 50 . what is order of reaction?

Grade:12th pass

1 Answers

Arun
5 years ago
t1/2 = 1 / A^(n-1)
 
now
 
0.1 / 0.4 = (50 /200)^(n-1)
 
¼ = ( ¼)^n-1
 
n – 1 = 1
 
n = 2
 
 
Regards
Arun (askIITians forum expert)